2 PbCl2 → Pb + PbCl4

Pb + Cl2 → PbCl2 ΔH = −223 kJ

PbCl2 + Cl2 → PbCl4 ΔH = −87 kJ


1
Expert's answer
2021-02-28T06:15:45-0500

Solution:

2PbCl2 → Pb + PbCl4

The following is known:

(1): Pb + Cl2 → PbCl2, ΔH = −223 kJ

(2): PbCl2 + Cl2 → PbCl4, ΔH = −87 kJ

Reaction 1 has to be flipped so that PbCl2 is a reactant. The value of ∆H is also changed from (-) to (+).

PbCl2 → Pb + Cl2, ΔH1 = +223 kJ

Reaction 2 does not need to be changed because PbCl4 is already a product.

PbCl2 + Cl2 → PbCl4, ΔH2 = −87 kJ

The two reactions are summed:

PbCl2 + PbCl2 + Cl2 → Pb + Cl2 + PbCl4

2PbCl2 → Pb + PbCl4, ΔHr

ΔHr = ΔH1 + ΔH2 = +223 kJ + (−87 kJ) = +136 kJ

ΔHr = 136 kJ.


Answer: ΔHr = 136 kJ.